# Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases

### 30 Jan 2024 v2

@@ 5.1 Setup
Datasets
We use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) for schema-level generalization and paraphrase settings. We also study generalization on SimpleQuestions-Balance (SQB) Wu et al. (2019) with unseen relations. WebQuestionsSP (WebQSP) Yih et al. (2016) is employed for cross-dataset transfer experiments as it is based on practical human-curated search logs. All experiments use S-expression as the logical form due to its clear and concise structure. Entity linking results are taken from TIARA Shu et al. (2022) for GrailQA and WebQSP, and ArcaneQA Gu and Su (2022) for GraphQuestions.
@@
We report the performance of comparison models from their papers. For the relation linking task on SQB, we use BERT Devlin et al. (2019) as the base model for GAIN. For KBQA tasks, we use the open-source advanced model TIARA Shu et al. (2022) as the base model for GAIN, due to its strong performance on zero-shot schema items [^5] Pangu Gu et al. (2022a) also uses entity linking results from TIARA.. TIARA is composed of multi-grained retrievers and a generator, with the retrievers providing KB contexts [^6] Entities, exemplary logical forms (ELF) and schema items are retrieved. for the generator. The term “TIARA+GAIN” represents a model (both the retrievers and the generator) that is first tuned using GAIN synthetic data and subsequently fine-tuned on a target dataset. For LLM evaluation, we use the latest gpt-3.5-turbo-0613777https://platform.openai.com/docs/models model, and the few-shot contexts are retrieved from the combination of GrailQA training set and synthetic dataset using the TIARA+GAIN retrievers.
@@
Following previous works, we use Exact Match (EM), F1, and Hits@1 to measure the performance of KBQA models. To evaluate adaptability to paraphrases (§3.2), we calculate the standard deviation (std) of EM/F1 for questions of each logical form template. As shown in Equation 1, suppose there are $n$ sets of paraphrases in the dataset, each set of paraphrases corresponds to a logical form template with $m$ natural language expressions, and the F1 score obtained by the KBQA model on the $j$-th question of the $i$-th set of paraphrases is $F1_{i,j}$. The metric $Std_{F1}$ first calculates the standard deviation of the F1 scores obtained by the model on the $m$ questions for each set of paraphrases and then calculates the average of the $n$ standard deviations. This metric is used to measure the robustness of the model to different representations of the same semantics, i.e., whether it can cope with diverse natural language expressions. A lower standard deviation indicates that the model is more adaptive to different expressions. $Std_{EM}$ is calculated in the same way.
#question #one-hop #two-hop #domain
127,329 78,668 48,661 759
#none #count #comparatives #superlatives
115,221 7,115 1,874 3,119
#class #relation #entity
5,078 12,942 46,645
#question #relation #subject #domain
162,557 7,349 108,804 673
@@ 5.2 Implementation Details
Our experiments are done on the machine with an NVIDIA A100 GPU and up to 504GB of RAM. We implement our models utilizing PyTorch Paszke et al. (2019) and Hugging Face [^8] https://huggingface.co/. TIARA+GAIN (T5-3B) takes about 100 hours to train the logical form generator on the synthetic dataset.
Training Question Generator
We fine-tune the T5-base model Raffel et al. (2020) to convert S-expression or triple to natural language questions. We set the beam size to 10, the learning rate to 3e-5, the number of epochs to 10, and the batch size to 8.
Training TIARA
The training of the TIARA model Shu et al. (2022) follows its original settings, including the setting of hyperparameters and the calculation of metrics. Note that Hits@1 on TIARA is obtained by randomly selecting one answer for each question 100 times. Both the schema retriever and generator of TIARA are pre-trained on synthetic data and then fine-tuned on KBQA datasets. Since GraphQuestions has no official training-valid split, we randomly take 200 questions from the original training set as the valid set.
Training the Relation Linking Model
We use the BERT-base-uncased model Devlin et al. (2019) to rank candidate relations for SQB, and the input form is the same as the schema retriever of TIARA. We set the learning rate to 3e-5, the batch size to 256, and the maximum number of epochs to 3 with early stopping.
5.3 Data Augmentation
The statistics of GAIN-synthetic datasets for both logical forms and triples are shown in Table 2 and 3999Details of synthetic data are shown in Appendix  B.. Note that theoretically, the sampling of the GAIN method is not limited to the scale of the synthetic data we use here.
@@ 6 Analysis
Table 7: Average standard deviation of EM and F1 scores for each set of paraphrases on the GrailQA valid set. ELF denotes exemplary logical form Shu et al. (2022).
Schema-Level Generalization
As shown in Tables 4 and 5, the models perform significantly better on i.i.d. than compositional and zero-shot generalization, with the zero-shot generalization being the most challenging. Besides, an increased number of model parameters, combined with richer data from GAIN, significantly enhance the generalization capabilities of T5 models. TIARA+GAIN achieves the highest EM scores, including that on zero-shot scenes. This demonstrates promising ideas for further improving LM generalization capabilities, i.e., the positive effect of synthetic data and parametric scales on training LMs. However, it is important to note that fine-tuned models consistently outperform few-shot learning models, regardless of whether the schema is seen or not. Given the training and inference costs of LLMs, their performance has yet to show any superiority in this semantic parsing task.
@@
Table 9: KBQA dataset statistics. Length denotes the average number of characters in the question. # of entities/relations denotes the average number of entities/relations contained in the logical form. Unseen Schema is the ratio of schema items in the dev/test set that do not appear in the training set. Unseen Question is the ratio of questions containing unseen schema.
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Valid Set EM F1 EM F1 EM F1 EM F1
BERT + Ranking (Gu et al., 2021) 51.0 58.4 58.6 66.1 40.9 48.1 51.8 59.2
TIARA ELF only Shu et al. (2022) 67.2 72.9 72.8 76.7 55.3 60.7 69.7 76.3
RnG-KBQA Ye et al. (2022) 71.4 76.8 86.7 89.0 61.7 68.9 68.8 74.7
DecAF (FiD-3B) Yu et al. (2022) - 81.4 - 89.7 - 80.1 - 78.4
TIARA (T5-base) Shu et al. (2022) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
Pangu (T5-3B) Gu et al. (2022a) 75.8 83.4 - - - - - -
TIARA + GAIN ELF only 67.4 73.6 72.7 77.5 54.7 60.5 70.3 77.3
TIARA + GAIN (T5-base) 77.1 83.5 89.0 91.9 68.6 75.5 75.4 83.2
TIARA + GAIN (T5-3B) 77.1 83.8 89.0 92.1 68.8 76.1 75.4 83.4
GPT-3.5-turbo (5-shot) 69.7 74.8 83.0 85.5 58.7 64.6 68.6 74.4
Paraphrase Adaptation
To evaluate the consistency of the model performance across varying expressions of the same semantics, we calculate the standard deviation (std) of EM or F1 for questions within each logical form template and subsequently calculate the average std for all templates in the dev/test set. A lower std signifies greater adaptability of the model to different expressions. For GrailQA, the std of EM and F1 decreases with the application of GAIN or an increase in model size. However, in the case of more challenging GraphQuestions, GAIN significantly improves the F1 but also results in a larger std, as shown in Table 5. It suggests that improving paraphrase adaptation seems more difficult when the baseline model (TIARA, T5-base, with only 37.9% F1) still struggles to address most of the dataset. Consequently, the performance gains observed on the KBQA benchmark may not necessarily reflect a deeper understanding of linguistic complexities, but they could simply render the model more sensitive to specific phrases.
Cross-Dataset Transfer
To emulate an unseen real-world scenario, we evaluate the performance of pre-trained models on the human-curated WebQSP dataset without fine-tuning, as depicted in Table 8. BERT+Ranking Gu et al. (2021) and TIARA+GAIN Shu et al. (2022) are trained on the large-scale GrailQA dataset. We compare these results to the state-of-the-art Pangu Gu et al. (2022a), which is fine-tuned on WebQSP and achieves an F1 score of 79.6%. Although we recognize that GAIN and large models offer few advantages, the performance of these pre-trained models without fine-tuning is considerably lower than Pangu’s. We attribute this to the significant differences between training and test data, as shown in Table 9. The question length, the difficulty of entity/relation linking [^10] Measured by literal similarity: https://anhaidgroup.github.io/py_stringmatching/v0.3.x/PartialRatio., and the proportion of unseen schema vary dramatically across KBQA datasets. These discrepancies arise from the dataset construction process: WebQSP is an annotation of search logs, whereas the remaining datasets are derived from graph search and crowdsourcing. To further enhance robustness in cross-dataset transfer, we believe that better data collection methods are required to obtain diverse and balanced training data. Additionally, the representation of the logical form increases the transfer difficulty, as the S-expression used in GrailQA dataset cannot express all queries in WebQSP.
LLM with In-context Learning
We evaluate the performance of GPT-3.5 with in-context learning on the GrailQA dataset. In the prompt, we provide the task description and the few-shot KB contexts. As illustrated in Table 10, when provided with contexts from the TIARA+GAIN retrievers, GPT-3.5 outperforms the ELF contexts but falls short compared to T5 generators. Among the GPT-3.5 predictions, 79.62% come directly from the substring of the corresponding prompts, achieving an average F1 score of 86.19% for this portion. However, the remaining predictions are not part of their prompts and are entirely new predictions generated by GPT-3.5, with an average F1 score of merely 30.29%. Although a baseline level is attained, these results suggest that GPT-3.5 cannot be accurately grounded to the KB environment when it does not directly utilize the retrievers’ contexts. It also means the severance of natural language pre-training and KB contexts for the LLM, and the faithfulness and controllability of grounding LLMs are not yet guaranteed under the current approach Gu et al. (2022a). To mitigate this problem, alternative paradigms should be explored, such as tool learning Schick et al. (2023) and multi-step planning Liu et al. (2023), which enables more refined access and control over environments.
@@ 7 Conclusion
We find that despite the voluminous training corpus available for LMs, incorporating data that adequately captures the intricate complexities of environments such as large-scale KBs is not yet fulfilled. This underscores the importance of focusing on the robustness challenges. Notably, LLMs sometimes blindly follow the provided prompt. It indicates that the existing methodologies for grounding LLMs are yet to prove their efficacy and superiority. Future research issues include collecting more balanced environment-specific corpora and improving the LLM learning paradigm. For the corpus collection problem, our experiments show some potential for data augmentation techniques.
@@
For question generation, the verbalization process of the GAIN method relies heavily on large-scale KBQA annotations. The generated questions may be similar to the training data, and overly complex logical forms (e.g., with three and more hops) are difficult to convert into natural language questions. Besides, synthetic data is less diverse and natural than human annotations, though it improves generalization performance.
@@
The proposed approach GAIN could be used on any KB for data augmentation. The Freebase Bollacker et al. (2008) used in this work is a KB that has been publicly released and manually reviewed. For uncensored KBs, if harmful information is collected, it could make synthetic data contain harmful information and make LMs generate harmful answers.
@@ Appendix B Details of Synthetic Data
The statistics for KBQA datasets including our synthetic dataset are shown in Table 16. To calculate the number of domains in Table 2, 3, we take the domain of each Freebase class from its first segment, except for classes starting with “base”, where we take the first two segments, e.g., domain “music” for the class “music.performance_venue”, and domain “base.plants” for the class “base.plants.plant”.

### 09 Feb 2024 v3

@@ 5 Experiments
5.1 Augmentation Approach
To ensure fair evaluation and fully harness the capabilities of LMs under our extensive and rigorous evaluation protocols, we suggest two strategies to counteract distribution shifts: data augmentation and retrieval augmentation.
5.1.1 Data Augmentation for LMs
Off-the-shelf datasets of limited size may make LM easily overfitted and not adaptable to large KBs. To address the problem that many domains in the KB are often not collected as training data, we propose a data augmentation method named Graph seArch and questIon generatioN (GAIN). Some data augmentation or question generation models Bi et al. (2020); Guo et al. (2022) are only evaluated by the qualify of generated sentences rather than evaluated by the QA task, but GAIN directly serves our KBQA evaluations. Besides, compared to the previous work Hu et al. (2019) that only considers generating questions for triples to help KBQA, GAIN applies to KBQA corresponding to both logical forms and triples. GAIN scales data volume and distribution through four steps: 1) Graph search: Sampling logical forms or triples from arbitrary domains in the KB without being restricted to any particular KBQA dataset. 2) Training question generator: learning to convert logical forms or triples into natural language questions on existing KBQA datasets. 3) Verbalization: Using the question generator from step 2 to verbalize sampled logical forms or triples from step 1, thus creating synthetic questions. 4) Training data expansion: Before fine-tuning any neural models on KBQA datasets, GAIN-synthetic data can be used to train these models or to expand the corpus of in-context samples for LLMs. That is, as a data augmentation method, GAIN is not a KBQA model, but it is used to augment a base KBQA model.
5.1.2 Retrieval Augmentation for LLMs
As the trade-off between cost and effectiveness, we experiment with the prevalent in-context learning paradigm but attempt to improve the quality of in-context samples. We use advanced retrieval methods based on smaller LMs as plug-ins to augment the LLM, similar to the SuperICL approach Xu et al. (2023). Specifically, our steps to generate an LLM prompt for each question include the following. 1) Given an input question, we retrieve $k$ questions ($k$-shot) with BM25 Robertson et al. (2009) from the corpus (the combination of KBQA training set and the GAIN-synthetic dataset). 2) The role of retrieval augmentation for KB environments has been shown by fine-tuned LMs Shu et al. (2022). To assist with grounding LLM, we retrieve KB contexts with off-the-shelf retrievers for $k$ samples and the input question. [^3] The prompt example is demonstrated in Appendix A.
@@ 5.2 Setup
Data
All experiments use S-expression Gu et al. (2021) as the logical form due to its clear and concise structure. Entity linking results are taken from TIARA Shu et al. (2022) for GrailQA and WebQSP, and ArcaneQA Gu and Su (2022) for GraphQuestions, because of their public availability and performance.
@@
Compared models are mainly selected from the leaderboard. [^4] https://dki-lab.github.io/GrailQA/ The performances are taken from their papers. For the relation linking task on SQB, we use BERT Devlin et al. (2019) as the base model for GAIN. For KBQA tasks, we use the open-source advanced model TIARA Shu et al. (2022) as the base model for GAIN, due to its strong performance on zero-shot schema items. [^5] Pangu Gu et al. (2022a) also uses entity linking results from TIARA. TIARA is composed of multi-grained retrievers and a generator, with the retrievers providing KB contexts [^6] Entities, exemplary logical forms, and schema items are retrieved. for the generator. The term “TIARA+GAIN” represents a model (both the retrievers and the generator) that is first tuned using GAIN synthetic data and subsequently fine-tuned on a target dataset. For LLM evaluation in the modal aspect, we use the gpt-3.5-turbo-0613777https://platform.openai.com/docs/models model, and the few-shot contexts are retrieved from the combination of GrailQA training set and synthetic dataset using the TIARA+GAIN retrievers.
@@
Following previous works, we use Exact Match (EM), F1, and Hits@1 to measure the performance of KBQA models. We also use the std of EM/F1 to measure the adaptability to paraphrases (§4.2).
@@ 5.3 Implementation Details
We use a machine with an NVIDIA A100 GPU and up to 504GB of RAM. Models are implemented by PyTorch Paszke et al. (2019) and Hugging Face. [^8] https://huggingface.co/ TIARA+GAIN (T5-3B) takes about 100 hours to train the logical form generator on the synthetic dataset.
Model Training
1) For question generation, we fine-tune the T5-base model Raffel et al. (2020) to convert S-expression or triple to natural language questions. We set the beam size to 10, the learning rate to 3e-5, the number of epochs to 10, and the batch size to 8. 2) The training of the TIARA model Shu et al. (2022) follows its original settings, including the setting of hyperparameters and the calculation of metrics. Note that Hits@1 on TIARA is obtained by randomly selecting one answer for each question 100 times. Both the schema retriever and generator of TIARA are pre-trained on synthetic data and then fine-tuned on KBQA datasets. Since GraphQuestions has no official training-valid split, we randomly take 200 questions from the original training set as the valid set. 3) We use BERT-base-uncased Devlin et al. (2019) to rank candidate relations for SQB, and the input form is the same as the schema retriever of TIARA. We set the learning rate to 3e-5, the batch size to 256, and the max number of epochs to 3 with early stopping.
Data Augmentation
The statistics of GAIN-synthetic datasets for both logical forms and triples are shown in Table 11 and 12.999Details of synthetic data are shown in Appendix B. Note that the sampling of the GAIN method is not limited to the scale of the synthetic data we use here.
@@ 6 Analysis
Table 5: Average standard deviation of EM and F1 scores for each set of paraphrases on the GrailQA valid set.
We report and analyze the experimental results in this section for each aspect.
6.1 Analysis of Environmental Aspect
Effectiveness of Synthesis and Scaling Up
As shown in Tables 2 and 3, the models perform significantly better on i.i.d. than compositional and zero-shot generalization, with the zero-shot partition being the most challenging. TIARA+GAIN (T5-base) improves 2.5 zero-shot F1 points compared to TIARA (T5-base). Besides, an increased number of model parameters, combined with richer data from GAIN, significantly enhance the generalization capabilities of T5 models. TIARA+GAIN (T5-3B) further improves 1.4 zero-shot F1 points compared to its T5-base version. TIARA+GAIN achieves the highest EM scores, including that on zero-shot scenes. It demonstrates promising ideas for further improving LM generalization capabilities, i.e., the positive effect of synthetic data and parametric scales on training LMs.
Fine-tuning Better Than Few-shot Learning
However, it is important to note that fine-tuned models consistently outperform few-shot learning models, regardless of whether the schema is seen or not. Given the training and inference costs of LLMs, their performance has yet to show any superiority in this task.
6.2 Analysis of Linguistic Aspect
Improvements Are Linguistic Biased
We calculate the standard deviation (std) of EM or F1 in the dev/test set, as shown in Equation 1. For GrailQA, the std of EM and F1 decreases with the application of GAIN or an increase in model size, i.e., F1/EM and the std of F1/EM are both better, as shown in Table 2 and 5. However, in the case of more challenging GraphQuestions, GAIN significantly improves the F1 by 8.3 points but also results in a larger std (0.170 compared to 0.157), as shown in Table 3. It suggests that improving paraphrase adaptation using GAIN is more difficult when the base model (TIARA, T5-base, with only 37.9% F1) still struggles to address most of the dataset. Consequently, the performance gains observed on the KBQA benchmark may not necessarily reflect a deeper understanding of linguistic complexities, but they could simply render the model more sensitive to specific phrases. Strategies for deeper decomposition and understanding Hu et al. (2021); Huang et al. (2023) of the questions may be needed to mitigate this challenge.
@@ 6.2 Analysis of Linguistic Aspect
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Valid Set EM F1 EM F1 EM F1 EM F1
BERT + Ranking (Gu et al., 2021) 51.0 58.4 58.6 66.1 40.9 48.1 51.8 59.2
TIARA ELF only Shu et al. (2022) 67.2 72.9 72.8 76.7 55.3 60.7 69.7 76.3
RnG-KBQA Ye et al. (2022) 71.4 76.8 86.7 89.0 61.7 68.9 68.8 74.7
DecAF (FiD-3B) Yu et al. (2022) - 81.4 - 89.7 - 80.1 - 78.4
TIARA (T5-base) Shu et al. (2022) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
Pangu (T5-3B) Gu et al. (2022a) 75.8 83.4 - - - - - -
TIARA + GAIN (T5-base) 77.1 83.5 89.0 91.9 68.6 75.5 75.4 83.2
TIARA + GAIN (T5-3B) 77.1 83.8 89.0 92.1 68.8 76.1 75.4 83.4
GPT-3.5-turbo (5-shot) 69.7 74.8 83.0 85.5 58.7 64.6 68.6 74.4
6.3 Analysis of Integrated Aspect
Hard Transfer Across Datasets
We evaluate the performance of pre-trained models on the human-curated WebQSP dataset without fine-tuning, as shown in Table 6. BERT+Ranking Gu et al. (2021) and TIARA+GAIN Shu et al. (2022) are trained on the large-scale GrailQA dataset. We compare these results to the state-of-the-art Pangu Gu et al. (2022a), which is fine-tuned on WebQSP and achieves an F1 score of 79.6% Although we recognize that GAIN and large models offer few advantages, the performance of these pre-trained models without fine-tuning is considerably lower than Pangu’s.
Causes from Data Collection
We attribute this to the significant differences between training and test data, as shown in Table 8. The question length, the difficulty of entity/relation linking [^10] Measured by literal similarity: https://anhaidgroup.github.io/py_stringmatching/v0.3.x/PartialRatio., and the proportion of unseen schema vary dramatically across KBQA datasets. These discrepancies arise from the dataset construction process: WebQSP is an annotation of search logs, whereas the remaining datasets are derived from graph search and crowdsourcing. To further enhance robustness in cross-dataset transfer, we believe that better data collection methods are required to obtain diverse and balanced training data. Additionally, the representation of the logical form increases the transfer difficulty, as the S-expression used in the GrailQA dataset cannot express all queries in WebQSP.
6.4 Analysis of Modal Aspect
Context Alone Is Insufficient
We evaluate the performance of GPT-3.5 using retrieved KB contexts (§5.1.2) and in-context learning on the GrailQA dataset. The prompts for the model include the task description and the few-shot KB contexts. As illustrated in Table 7, when provided with contexts from the TIARA+GAIN retrievers, GPT-3.5 outperforms two compared models but falls short compared to TIARA+GAIN. Among the GPT-3.5 predictions, 79.62% come directly from the substring of the corresponding prompts, achieving an average F1 score of 86.19% for this portion. However, the remaining predictions are not part of their prompts and are entirely new predictions generated by GPT-3.5, with an average F1 score of merely 30.29%. Although a baseline level is attained, these results suggest that GPT-3.5 cannot be accurately grounded to the KB environment when it does not copy the retrievers’ contexts. It also shows the modal severance of natural language pre-training and KB contexts for the LLM. LLMs are not able to utilize KB contexts as proficiently as they understand natural language, and the faithfulness and controllability of grounded LLMs are not yet guaranteed under the current approach Gu et al. (2022a). To mitigate this problem, alternative paradigms should be explored, such as tool learning Schick et al. (2023) and multi-step planning Liu et al. (2023) with Chain-of-Thought Wei et al. (2022), which enables more refined access and control over environments and reduces modal differences during the reasoning process.
@@ 6.4 Analysis of Modal Aspect
Table 8: KBQA dataset statistics. Length denotes the average number of question characters. # of entities/relations denotes the average number of entities/relations in the logical form. Unseen Schema is the ratio of unseen schema items in the dev/test set. Unseen Question is the ratio of questions containing unseen schema.
@@ 7 Conclusion
Despite the recent progress of LM-driven models, robustness challenges posed by distribution shifts for the KBQA task are rarely discussed. Our analyses call for further research into better evaluation protocols for grounding LMs to KBs and enhancing the robustness of environmental, linguistic, and modal aspects. Notably, the experiments reveal that LLMs sometimes simply copy the provided prompt. It indicates that the existing methodologies for grounding LLMs are yet to prove their efficacy and superiority. Future research issues include collecting more balanced environment-specific corpora and improving the LLM learning paradigms. For the corpora, our experiments show that the data augmentation techniques deserve further research.
@@
1) For question generation, the verbalization process of the GAIN method relies heavily on large-scale KBQA annotations. The training data influence the style of generated questions, and overly complex logical forms (e.g., with three or more hops) are difficult to convert into natural language questions. Besides, synthetic data is less diverse and natural than human annotations, though it improves generalization performance. 2) Multilingual KBQA is a problem that lacks attention in the KBQA research and is also a linguistic-aspect challenge. However, since most KBQA datasets are based on English and do not discuss other languages at all, this paper leaves the evaluation of this problem for future work.
@@
The proposed data augmentation method GAIN could be used on any KB. The Freebase Bollacker et al. (2008) used in this work is a KB that has been publicly released and manually reviewed. For uncensored KBs, if harmful information is collected, it could make synthetic data contain harmful information and make LMs generate harmful answers.
Acknowledgments
We extend our sincere gratitude to all the anonymous reviewers for their insightful suggestions. Special thanks to Xiao Xu, Xiang Huang, and Sitao Cheng for their valuable discussions on this paper. We appreciate Yu Gu’s efforts in evaluating our submissions on the GrailQA benchmark test set.
@@ Appendix A Example of LLM Prompt
#question #one-hop #two-hop #domain
127,329 78,668 48,661 759
#none #count #comparatives #superlatives
115,221 7,115 1,874 3,119
#class #relation #entity
5,078 12,942 46,645
#question #relation #subject #domain
162,557 7,349 108,804 673
@@ Appendix B Details of Synthetic Data
The statistics for KBQA datasets, including our synthetic dataset, are shown in Table 16. To calculate the number of domains in Table 11, 12, we take the domain of each Freebase class from its first segment, except for classes starting with “base”, where we take the first two segments, e.g., domain “music” for the class “music.performance_venue”, and domain “base.plants” for the class “base.plants.plant”.